# Towards Adapting Open-Source Large Language Models for Expert-Level Clinical Note Generation

### v5

@@ Appendix D Additional Results
D.1 Qualitative Analysis of Model Outputs
@@
D.2 Physician Reader Study
Two internal medicine physicians and one family medicine physician, in a blinded review, evaluated notes authored by physicians, LLaMA-Clinic, and Gemini Pro based on three criteria: real-world readiness, completeness, and accuracy (Figure 4A). The median word counts and interquartile ranges (IQR) for notes authored by physicians, LLaMA-Clinic, and Gemini Pro were 118 (IQR: 94-150), 128 (IQR: 108-145), and 128 (IQR: 100-164), respectively. No statistically significant differences in word counts were observed among the three groups (Kruskal-Wallis H test: p = 0.292). We assessed inter-rater reliability (IRR) utilizing Gwet’s AC2 statistics. The AC2 scores for the three metrics ranged from 0.80 to 0.83, signifying a high degree of agreement among reviewers.
@@ D.2 Physician Reader Study
D.3 Cost Analysis for Model Development and Inference
@@ Appendix E Additional Discussion
Continued pretraining of an LLM using a domain-specific corpus is recognized for enhancing performance on downstream tasks wu2023bloomberggpt. This phase is considered a knowledge injection process, given that LLMs acquire the vast majority of their knowledge during the pre-training phase ovadia2023fine; zhou2023lima. Several clinical LLMs that underwent continued pretraining with medical corpora, such as PubMed literature, combined with SFT have shown significant improvements in medical knowledge benchmarks chen2023meditron; wu2023pmcllama; luo2023biomedgpt.
However, a critical distinction exists between tasks focused on medical knowledge (e.g., answering USMLE questions) and those aimed at clinical note generation. Notably, GatorTronGPT peng2023study, the only LLM trained from scratch using EMR data from real patients to date, performed lower in both MedQA and PubMedQA compared to other clinical LLMs chen2023meditron; wu2023pmcllama; kweon2023publicly. This outcome indicates that EMR data alone may lack comprehensive medical knowledge. For the task of note generation, we hypothesized that continued pretraining with clinical notes could offer benefits by introducing greater lexical variance, unique semantic patterns, and diverse formatting similar to prior work lehman2023we. Our experiments did not conclusively demonstrate the anticipated benefits of continued pretraining, as the out-of-the-box LLaMA-2 model without continued pretraining achieved the highest ROUGE-1 scores after SFT and RLAIF (Appendix Table 2). We opted to proceed with the continued pretrained model for RLHF due to subtle peculiarities observed upon manual inspection. Given the significant time and computational resources required for continued pretraining, its utility, particularly with clinical notes, merits further exploration in future work.
@@ E.2 Data Selection for Continued Pretraining
Another potential factor in the less impressive improvement from continued pretraining may be attributed to the variance in data distribution between discharge summaries and outpatient notes. To address this, we performed experiments focused on a condensed version of discharge summaries, hypothesizing that the “brief hospital course” section would contain data of higher quality than the complete discharge summary. Indeed, models trained on the “brief hospital course” section outperformed those trained on the full summaries (Table 1). Interestingly, we observe that during the pretraining stage, models trained on full summaries achieved lower training losses (see Figure 3). However, this did not lead to better performance in the downstream task. We speculate that the structured nature of the full discharge summaries, which include sections such as laboratory results and medication lists, presents more straightforward learning targets for the model. This allows it to achieve lower training losses, which do not necessarily translate into improved task performance. This observation underscores the necessity for thorough analysis of the data used for pretraining.
@@ 2. Hypertension
Appendix Table 12: Pricing Information for Inference Cost Estimation. We sourced pricing information from the websites of Google AI GoogleAIPricing2024, OpenAI OpenAIAPIPricing2024, and Fireworks.ai FireworksAI2024 in May 2024 for on-demand API calls. The price of GPT-3.5 Turbo is based on the model gpt-3.5-turbo-1106. Fireworks.ai charges the same price for both input and output tokens.

### v6

@@ Appendix D Additional Results
D.1 Performance of LLM-as-a-Judge in Predicting Physician Preference
We evaluated Gemini 1.0 as an LLM-as-a-judge on our RLHF dataset. We observed a low accuracy of 43.8% in predicting physician preferences (from three candidate notes), underscoring the challenges LLMs encounter in predicting expert preferences in complex domains such as medicine.
D.2 Qualitative Analysis of Model Outputs
@@
D.3 Physician Reader Study
Four internal medicine physicians and one family medicine physician, in a blinded review, evaluated notes authored by physicians, LLaMA-Clinic, and Gemini Pro based on three criteria: real-world readiness, completeness, and accuracy (Figure 4A). The median word counts and interquartile ranges (IQR) for notes authored by physicians, LLaMA-Clinic, and Gemini Pro were 118 (IQR: 94-150), 128 (IQR: 108-145), and 128 (IQR: 100-164), respectively. No statistically significant differences in word counts were observed among the three groups (Kruskal-Wallis H test: p = 0.292). We assessed inter-rater reliability (IRR) utilizing Gwet’s AC2 statistics. The AC2 scores for the three metrics ranged from 0.80 to 0.82, signifying a high degree of agreement among reviewers.
@@ D.3 Physician Reader Study
D.4 Cost Analysis for Model Development and Inference
@@ Appendix E Additional Discussion
Continued pretraining of an LLM using a domain-specific corpus is recognized for enhancing performance on downstream tasks Wu et al. (2023b). This phase is considered a knowledge injection process, given that LLMs acquire the vast majority of their knowledge during the pretraining phase Ovadia et al. (2023); Zhou et al. (2023a). Several clinical LLMs that underwent continued pretraining with medical corpora, such as PubMed literature, combined with SFT have shown significant improvements in medical knowledge benchmarks Chen et al. (2023); Wu et al. (2023a); Luo et al. (2023).
However, a critical distinction exists between tasks focused on medical knowledge (e.g., answering USMLE questions) and those aimed at clinical note generation. Notably, GatorTronGPT Peng et al. (2023), the only LLM trained from scratch using EMR data from real patients to date, performed lower in both MedQA and PubMedQA compared to other clinical LLMs Chen et al. (2023); Wu et al. (2023a); Kweon et al. (2023); Wang et al. (2024b). This outcome indicates that EMR data alone may lack comprehensive medical knowledge. For the task of note generation, we hypothesized that continued pretraining with clinical notes could offer benefits by introducing greater lexical variance, unique semantic patterns, and diverse formatting similar to prior work Lehman et al. (2023). Our experiments did not conclusively demonstrate the anticipated benefits of continued pretraining. Interestingly, the LLaMA-2 model without continued pretraining achieved the highest ROUGE-1 scores after SFT and RLAIF in our early experiments (see Appendix Table 2). We opted to proceed with the continued pretrained model for RLHF due to subtle peculiarities observed upon manual inspection. Given the significant time and computational resources required for continued pretraining, its utility, particularly with clinical notes, merits further exploration in future work.
@@ E.2 Data Selection for Continued Pretraining
Another potential factor in the less impressive improvement from continued pretraining may be attributed to the variance in data distribution between discharge summaries and outpatient notes. To address this, we performed experiments focused on a condensed version of discharge summaries (Discharge-short), hypothesizing that the “brief hospital course” section would contain data of higher quality than the complete discharge summary. Indeed, models trained on the Discharge-short outperformed those trained on the full summaries (Table 1). Interestingly, we observe that during the pretraining stage, models trained on full summaries achieved lower training losses (see Figure 3). However, this did not lead to better performance in the downstream task. We speculate that the structured nature of the full discharge summaries, which include sections such as laboratory results and medication lists, presents more straightforward learning targets for the model. This allows it to achieve lower training losses, which do not necessarily translate into improved task performance. This observation underscores the necessity for thorough analysis of the data used for pretraining.
@@ 2. Hypertension
Appendix Table 11: Pricing Information for Inference Cost Estimation. We sourced pricing information from the websites of Google AI Goo (2024), OpenAI Ope (2024), and Fireworks.ai Fir (2024) in May 2024 for on-demand API calls. The price of GPT-3.5 Turbo is based on the model gpt-3.5-turbo-1106. Fireworks.ai charges the same price for both input and output tokens.